OpenStudy (anonymous):

Ill medal if u help me and fan as well What is 5 * 1 3/5=

OpenStudy (anonymous):

@ravege

OpenStudy (anonymous):

@YoungStudier

OpenStudy (just_one_last_goodbye):

@Thaqib need help?

OpenStudy (anonymous):

yes please?

OpenStudy (anonymous):

\[5 \left(1+\frac{3}{5}\right)=8 \]

OpenStudy (just_one_last_goodbye):

aww ;-;

OpenStudy (anonymous):

thanks i knew its was 8 thx

OpenStudy (anonymous):

dont worry u can help me with another one and u get a medel 2

OpenStudy (mrnood):

so why ask - if oyu KNEW?

OpenStudy (anonymous):

What is 3 3/8 * 4 1/4= ? Express the answer in simplest form.

OpenStudy (anonymous):

btw the answer should be a whole number

OpenStudy (mrnood):

convert both terms to improper fractions then multiply both the numerators multiply both denominators simplify the result

OpenStudy (anonymous):

ok so 24/8 and 13/3?

OpenStudy (anonymous):

are those the improper fractions

OpenStudy (mrnood):

no - it's not 24/8 - you didnt ad the 3/8

OpenStudy (anonymous):

so 11

OpenStudy (mrnood):

and the second one is ?/4 stop rushing 11 is not correct either

OpenStudy (anonymous):

What is 3 3/8 * 4 1/4= ? Express the answer in simplest form.

OpenStudy (mrnood):